Activity in Limit Theorems



Directions: Assume the following.



1. lim f(x) = 3/4


x→c



2. lim g(x) = 12


x→c



3. lim h(x) = -3


x→c






1
Expert's answer
2022-03-15T11:59:01-0400

Answer


Given that


"\lim_{x\rightarrow c}f(x)=\frac{3}{4}" 


"\lim_{x\rightarrow c}g(x)=12" 


"\lim_{x\rightarrow c}h(x)=-3"


We can use these values in the following examples to understand some of the limit theorems.

limit theorems.


Ex 1. "\lim_{x\rightarrow c}f(x)\pm\lim_{x\rightarrow c}g(x)"


To find "\lim_{x\rightarrow c}f(x)\pm\lim_{x\rightarrow c}g(x)" , we use the given values,


"\lim_{x\rightarrow c}f(x)=\frac{3}{4}" and "\lim_{x\rightarrow c}g(x)=12"


Therefore,


"=\frac{3}{4}\pm(12)"


"=\frac{3}{4}+(12)=\frac{51}{4}" and "=\frac{3}{4}-(12)=\frac{-45}{4}"




Ex 2. "\lim_{x\rightarrow c}[(2.f(x))+\sqrt{12.g(x)}]"


To find "\lim_{x\rightarrow c}[(2.f(x))+\sqrt{12.g(x)}]" , we use the given values,


"\lim_{x\rightarrow c}[(2.f(x))+\sqrt{12.g(x)}]\\=[(2.\lim_{x\rightarrow c}f(x))+\sqrt{12.\lim_{x\rightarrow c}g(x)}]"


"=[(2.\frac{4}{3})+\sqrt{(12)(12)}]"


"=\frac{8}{3}+12=\frac{44}{3}"


"\lim_{x\rightarrow c}f(x)=\frac{3}{4}" and "\lim_{x\rightarrow c}g(x)=12"


Therefore,


"=2(\frac{3}{4})+(12)=\frac{27}{2}"



Ex 3. "lim_{x\rightarrow c}4\frac{f(x)-g(x)}{2h(x)}"


"=\frac{\lim_{x\rightarrow c}4f(x)-\lim_{x\rightarrow c}g(x)}{\lim_{x\rightarrow c}2h(x)}"


"=\frac{4\lim_{x\rightarrow c}f(x)-\lim_{x\rightarrow c}g(x)}{2\lim_{x\rightarrow c}h(x)}"


"=\frac{4(\frac{3}{4})-(12)}{(2)(-3)}"


"=\frac{3}{2}"



Ex 4. "lim_{x\rightarrow c}[h(x)-4f(x)]"


"=lim_{x\rightarrow c}h(x)-4lim_{x\rightarrow c}f(x)"

"=(-3)-4(\frac{3}{4})"


"=-3-3"


"=-6"